# Navigating Rifts in Human-LLM Grounding: Study and Benchmark

### v1

@@ Abstract
Language models excel at following instructions but often struggle with the collaborative aspects of conversation that humans naturally employ. This limitation in grounding—the process by which conversation participants establish mutual understanding—can lead to outcomes ranging from frustrated users to serious consequences in high-stakes scenarios. To systematically study grounding challenges in human-LLM interactions, we analyze logs from three human-assistant datasets: WildChat, MultiWOZ, and Bing Chat. We develop a taxonomy of grounding acts and build models to annotate and forecast grounding behavior. Our findings reveal significant differences in human-human and human-LLM grounding: LLMs were three times less likely to initiate clarification and sixteen times less likely to provide follow-up requests than humans. Additionally, early grounding failures predicted later interaction breakdowns. Building on these insights, we introduce Rifts: a benchmark derived from publicly available LLM interaction data containing situations where LLMs fail to initiate grounding. We note that current frontier models perform poorly on Rifts, highlighting the need to reconsider how we train and prompt LLMs for human interaction. To this end, we develop a preliminary intervention that mitigates grounding failures.
@@ 1 Introduction
Language models used for conversational interaction are trained primarily to follow instructions Ouyang et al. (2022). But effective dialogue requires more than just instruction-following. Participants in conversation work together in a collaborative process, resolving ambiguities as they exchange ideas and achieve shared objectives. They cultivate common ground111We use grounding to refer to Clark’s formulation of the language, gestures, and signaling that participants in a conversation employ to establish and maintain an effective dialog with shared understanding among actors Clark (1996)—in distinction to “grounding” in multimodal grounding. through grounding processes: communicative behaviors aimed at establishing and confirming mutual comprehension Clark (1996). Grounding mechanisms guide how individuals anticipate and react to a conversation partner’s contributions (Clark and Schaefer, 1989). Speakers implement grounding through dialogue acts. They work to confirm or clarify assumptions, and ask follow-up questions. When grounding breaks down, participants employ repair strategies to resolve potential miscommunications.
@@
In contrast, LLMs employed in conversational systems generate task-centric content with minimal grounding actions Shaikh et al. (2024). Failing to ground with users can be costly, with outcomes ranging from the common situation of frustrated users to that of serious consequences in high-stakes situations (Figure 1). The process by which humans build common ground offers a powerful metaphor for understanding human-computer interaction Brennan (2014). Ideal human-LLM grounding should allow for both human and machine initiatives, aimed at detecting misunderstandings and achieving mutual understanding Horvitz (1999).
@@
We start by assessing the current state of grounding in human-LLM interactions, analyzing grounding behavior through real-world interaction logs. To examine human-LLM grounding, we first characterize interactions through a set of validated dialogue acts (§3). In synthesizing these grounding acts, we build on prior work in conversational analysis and dialog systems. We cover acts that communicate progress (acknowledgment, follow-up, etc.) and difficulty in grounding (repair, clarification, etc.), initiated by either a human or LLM.
@@
Using a set of grounding acts, we construct models of human-LLM grounding, which we use to annotate and forecast grounding in conversation (§4-5). We construct and validate an LLM-based annotator, which enables efficient annotation of logs from publicly available interactions with ChatGPT (WildChat), data from a widely used commercial LLM service (Bing Chat), and Wizard-of-Oz-ed interactions with a human roleplaying as an AI (MultiWOZ). We additionally develop a grounding forecaster that predicts the presence of grounding acts in future turns. The forecaster enables a dialog intervention that can proactively prevent grounding difficulties in forthcoming turns.
@@
Our analysis of human-LLM interaction data reveals significant asymmetries in initiating grounding: people are three times more likely to clarify and sixteen times more likely to issue follow-up requests compared to LLMs. In addition, we find that grounding failures occurring early in dialog cascade into higher likelihoods of downstream failure. Overall, there is substantial room for improvement in human-LLM interaction via addressing deficits in conversational grounding.
@@ 2 Related Work
Disambiguating questions like “Did you see (the man (with the telescope))?” requires establishing common ground. While NLP benchmarks address ambiguity Min et al. (2020); Tamkin et al. (2022), they focus on well-defined cases (e.g., reference ambiguity). However, people engage LLMs for open-ended tasks—creative writing, cover letters—where correct answers aren’t predefined, making common ground crucial. Using naturalistic human-LLM interactions, we identify challenges in building common ground. To operationalize it Clark and Schaefer (1989); Clark (1996); Stalnaker (2002), we synthesize dialogue acts based on subdialogues Litman (1985); Litman and Allen (1987) and conversation structure Jefferson (1972).
Grounding in Dialog Systems
NLP systems from ELIZA onward have been designed to initiate some form of conversational grounding Weizenbaum (1966); Purver (2004); Li et al. (2023); Paranjape and Manning (2021). Decision-theoretic models have helped systems manage uncertainty about user goals Horvitz and Paek (2000a, b); Paek and Horvitz (2003), while multimodal approaches consider language and visual cues Pejsa et al. (2014). Human-LLM grounding is crucial for tasks including goal-coordination Bara et al. (2021); Mohanty et al. (2023); Fried et al. (2022); Li and Boyer (2015), planning Chu-Carroll and Carberry (1998); Lochbaum (1998), games Madureira and Schlangen (2023b); Shaikh et al. (2023), data retrieval Lu et al. (2023), improvisation Cho and May (2020), and design Vaithilingam et al. (2024). AI-initiated grounding improves conversation quality Zhou et al. (2022), enables human-AI collaboration Lin et al. (2023), and encorages humans to reflect on LLM outputs Park and Kulkarni (2023). Our work extends this by introducing methods and a benchmark for studying grounding in real-world human-LLM dialogue.
@@
Research has explored using machine-learned models to predict and mitigate grounding failures in spoken dialog systems. By forecasting potential failures, models can guide proactive interventions. For example, in a call-routing system, predictions of downstream grounding issues were used to trigger early transfers to human operators, reducing user frustration and disengagement, like pressing keys to access human assistants or abandoning calls Horvitz and Paek (2007).
@@
| Continuing the conversation with the next relevant turn. | ||
| Note: all other successful acts are a subset of Continue. | ||
@@ 3 Human-LLM Grounding Acts
To measure grounding between people and LLMs, we curate a set of dialogue acts that serve as proxies for grounding. We outline our selected acts and discuss prior work motivating each act. Our typology builds on prior work in conversational grounding: Clark and Schaefer (1989) and Traum and Hinkelman (1992) outline a hierarchy of actions—including discourse acts—that are used by humans to ground with one another. Recent work has revisited conversational grounding in the context of LLMs Shaikh et al. (2024); Schneider et al. (2024), focusing on a subset of acts generated mainly by people (e.g, following-up, acknowledging understanding, and clarifying).
In contrast, we consider acts generated by both LLMs and people. Our selected acts serve as signals for effective grounding; we segment acts across communicated grounding outcomes. We focus on the construction of common ground, a potential grounding failure, or acknowledging . Using our typology, we can measure grounding outcomes with observable dialog acts during human-LLM interaction (illustrative examples in Table 1).
@@
Addressing acts explicitly signal a potential misunderstanding. Here, participants are addressing this failure in an interaction.
occur when a participant repeats or restates their query because of a failure to ground. An utterance is a reformulation if the succeeding utterance from the same participant is semantically equivalent to the original. Reformulations are prevalent in search engine and information retrieval domains Lau and Horvitz (1999).
occur when users reset a conversation to improve understanding and achieve a successful dialog. They often follow significant misunderstandings, whether in the initial response or across multiple exchanges. Users may restart due to LLMs misinterpreting intent, ambiguity, sensitivity, or irrelevant context Shi et al. (2023), akin to search query retries after irrelevant results Lau and Horvitz (1999). Research on restarts includes user decisions to suspend a dialog and seek alternate solutions, such as transferring from AI-based systems to human operators before frustration escalates Horvitz and Paek (2007). Restarts are classified when an initial instruction is repaired or reformulated within 30 minutes Radlinski and Joachims (2005); Downey et al. (2007).
@@
Grounding may also be uncertain between participants. Rather than clearly indicating success or failure, ambiguous acts represent strategies that participants use to—potentially inefficiently—prevent potential misunderstandings.
occur when a participant disambiguates an utterance from another participant; or when a participant proactively “clears up” misunderstandings. Clarifications often occur when a task is perceived as ambiguous (e.g. What did you mean by that?). Ginzburg and Cooper (2001); Purver et al. (2003b, a); Healey et al. (2011, 2003); Purver (2004); Stoyanchev et al. (2013); Madureira and Schlangen (2023a); Rahmani et al. (2023).
include more than what another participant reasonably asked for. Unlike Next Turn, which provides only expected information, overresponses also anticipate potential follow-ups. They often appear as overly verbose LLM-generated answers, a behavior contemporary reward models are criticized for encouraging by favoring longer responses Singhal et al. (2023).
@@
Finally, advancing acts communicate that a participant understood past utterances from another participant in an interaction.
refer to the next conversational move made by a listener that is expected, given the prior turn(s) in a dialogue. Examples of relevant next turns include directly answering a question, expressing an opinion (agreeing or disagreeing), or apologizing. If no misunderstanding has occurred, a listener has moved on to the next relevant turn by default. Levinson et al. (1983); Sacks et al. (1978); Schiffrin (1987). Note that advancing grounding will initiate the Next Relevant Turn. We focus on two: follow-ups & acknowledgments.
elaborate on a prior utterance in an interaction. Unlike clarifications—which disambiguate or clarify—follow-ups seek additional information. Because follow-ups build on a prior utterance, they implicitly signal understanding of past utterances. Davis (1982); Graesser et al. (1995); Traum and Hinkelman (1992); Bunt et al. (2017).
explicitly signal understanding. These requests manifest either through explicit dialogue (e.g. I see; I understand; O.K.) or by repeating portions of another participant’s utterance (e.g. I can help you [write a story].) Unlike Reformulation, where a listener repeats to address failure, acknowledgment occurs when a speaker repeats to demonstrate understanding. Schegloff (1982); Sacks et al. (1978); Schiffrin (1987); Clark and Schaefer (1989); Cho and May (2020)
@@ 4 Data
To analyze collaborative grounding with LLM-based assistants, we draw from three English datasets consisting of dialogs between a human and an assistant. WildChat is a real-world Human-AI interaction dataset with interaction between people and several OpenAI models Zhao et al. (2024). User data was collected with consent, in exchange for free access to the models. We use the non-toxic version of WildChat, filter conversations to be in English, and sample one conversation from each user. Bing Chat, similar to WildChat, was collected from a large chat-based service used by millions of users, powered by OpenAI LLMs. Finally, MultiWOZ is a crowdsourced dataset of dialogue-based interaction with an assistant Budzianowski et al. (2018); Ramadan et al. (2018); Eric et al. (2019); Zang et al. (2020). In contrast to WildChat and Bing Chat, MultiWOZ contains human-human dialogs, with one human playing a “wizard-of-oz” role as the assistant. We use MultiWOZ 2.2 Zang et al. (2020), examining collaborative grounding acts on the validation and test splits for a subset of the tasks. While we observe similarities in terms of the tasks posed by humans across the three datasets, differences do exist, which makes direct comparisons difficult. In Appendix A.1, we outline the number of dialogs and messages in each dataset.
@@ 5.1 Classifying Human-LLM Grounding Acts
Method
We employ GPT-4o-mini to annotate grounding acts across a subset of our datasets. On subsets of our data, we observed nearly identical results using GPT-4o compared to GPT-4o-mini. Thus, we employ GPT-4o-mini for its affordability to allow for efficiency and reproducibility. Following Shaikh et al. (2024), we first encode our typology in a prompt (Appendix E.1) and label each turn in the conversations. To validate the accuracy of the approach, three authors annotated 10 dialogues (total of 108 messages) from each dataset and reached an average Cohen Kappa agreement of 0.71 across the datasets. Ties were broken through discussion, converging on a final selection of the majority label. Using this as a withheld test set, we find that Macro F-1 scores are reasonably high across datasets (0.75). A full table of results across labels is also in Appendix E.1.
Results
@@
Before a human ends up addressing grounding, an ideal LLM assistant would have proactively engaged in clarification. However, we find that LLMs clarify at significantly lower rates (, t-test) compared to humans repairing. In fact, the opposite occurs: people clarify LLM outputs ( ) 3 times as much as LLMs clarify user instructions ( ; Fig. 3). In contrast, humans and wizard-of-oz-ed assistants clarify at similar rates (MultiWOZ; human user v.s. human assistant).
@@
Early grounding patterns also have rippling effects: we find evidence of compounding advancing and addressing patterns as a conversation progresses (Fig. 5). Let denote the message at turn , which can be an advancing grounding act or addressing where and denote the sets of advancing and addressing grounding acts. The likelihood of an addressing utterance grounding appearing in the first turn on WildChat, and on Bing Chat. We observe a compounding of addressing grounding in human-assistant conversations. If we assume that the turn before was addressing, we find that the likelihood of another failed interaction triples: for Wildchat and for Bing Chat. A similar effect appears for advancing grounding: early signs of effective conversational grounding similarly snowballs, with on Wildchat and on Bing Chat.
5.2 Forecasting a User’s Grounding Acts
The GPT-based annotator (from §5.1) annotates conversations with grounding acts in a post-hoc fashion; we only identify opportunities for grounding after they happen. In contrast, curating tasks where grounding failures emerge requires predicting the likelihood of future grounding patterns. We focus on WildChat as it is the only publicly availible human-LLM interaction dataset in our analysis.
| Category | Prompt |

### v2

@@ Abstract
Language models excel at following instructions but often struggle with the collaborative aspects of conversation that humans naturally employ. This limitation in grounding—the process by which conversation participants establish mutual understanding—can lead to outcomes ranging from frustrated users to serious consequences in high-stakes scenarios. To systematically study grounding challenges in human-LLM interactions, we analyze logs from three human-assistant datasets: WildChat, MultiWOZ, and Bing Chat. We develop a taxonomy of grounding acts and build models to annotate and forecast grounding behavior. Our findings reveal significant differences in human-human and human-LLM grounding: LLMs were three times less likely to initiate clarification and sixteen times less likely to provide follow-up requests than humans. Additionally, we find that early grounding failures predict later interaction breakdowns. Building on these insights, we introduce Rifts, a benchmark derived from publicly available LLM interaction data containing situations where LLMs fail to initiate grounding. We note that current frontier models perform poorly on Rifts, highlighting the need to reconsider how we train and prompt LLMs for human interaction. To this end, we develop a preliminary intervention aimed at mitigating grounding failures.
@@ 1 Introduction
Language models used for conversational interaction are trained primarily to follow instructions Ouyang et al. (2022). But effective dialogue requires more than just instruction-following. Participants in conversation work together in a collaborative process, resolving ambiguities as they exchange ideas and achieve shared objectives. They cultivate common ground111We use grounding to refer to Clark’s formulation of the language, gestures, and signaling that participants in a conversation employ to establish and maintain an effective dialogue with shared understanding among actors Clark (1996). through grounding processes: communicative behaviors aimed at establishing and confirming mutual comprehension Clark (1996). Grounding mechanisms guide how individuals anticipate and react to a conversation partner’s contributions (Clark and Schaefer, 1989). Speakers implement grounding through dialogue acts. They work to confirm or clarify assumptions, and ask follow-up questions. When grounding breaks down, participants employ repair strategies to resolve potential miscommunications.
@@
In contrast, LLMs employed in conversational systems generate task-centric content with minimal grounding actions Shaikh et al. (2024). Failing to ground with users can be costly, with outcomes ranging from the common situation of frustrated users to that of serious consequences in high-stakes situations (Figure 1). The process by which humans build common ground offers a useful lens through which tounderstand human-computer interaction Brennan (2014). Ideal human-LLM grounding should allow for both human and machine initiatives, aimed at detecting misunderstandings and achieving mutual understanding Horvitz (1999).
@@
We start by assessing the current state of grounding in human-LLM interactions, analyzing grounding behavior through real-world interaction logs. To examine human-LLM grounding, we first characterize interactions through a set of validated dialogue acts (§3). In synthesizing these grounding acts, we build on prior work in conversational analysis and dialogue systems. We cover acts that communicate progress (acknowledgment, follow-up, etc.) and difficulty in grounding (repair, clarification, etc.), initiated by either a human or LLM.
@@
Using a set of grounding acts, we construct models of human-LLM grounding, which we use to annotate and to forecastgrounding in conversation (§4-5). We construct and validate an LLM-based annotator, which enables efficient annotation of logs from publicly available interactions with ChatGPT (WildChat), data from a widely used commercial LLM service (Bing Chat), and Wizard-of-Oz-ed interactions with a human roleplaying as an AI (MultiWOZ). We additionally develop a grounding forecaster that predicts the presence of grounding acts in future turns. The forecaster enables a dialogue intervention that can proactively prevent grounding difficulties in forthcoming turns.
@@
Our analysis of human-LLM interaction data reveals significant asymmetries in initiating grounding: people are three times more likely to clarify and sixteen times more likely to issue follow-up requests compared to LLMs. In addition, we find that grounding failures occurring early in dialogue cascade into higher likelihoods of downstream failure. Overall, there is substantial room for improvement in human-LLM interaction via addressing deficits in conversational grounding.
@@ 2 Related Work
Disambiguating questions like “Did you see (the man (with the telescope))?” requires establishing common ground. While NLP benchmarks address ambiguity Min et al. (2020); Tamkin et al. (2022), they focus on well-defined cases (e.g., reference ambiguity). However, people engage LLMs for open-ended tasks, e.g., creative writing and cover letters—where correct answers aren’t predefined, making common ground crucial. Using naturalistic human-LLM interactions, we identify challenges in building common ground. To operationalize it Clark and Schaefer (1989); Clark (1996); Stalnaker (2002), we synthesize dialogue acts based on subdialogues Litman (1985); Litman and Allen (1987) and conversation structure Jefferson (1972).
Grounding in Dialogue Systems
Numerous NLP systems from ELIZA onward have been designed to initiate some form of conversational grounding Weizenbaum (1966); Purver (2004); Li et al. (2023); Paranjape and Manning (2021). Decision-theoretic models have helped systems manage uncertainty about user goals Horvitz and Paek (2000a, b); Paek and Horvitz (2003), while multimodal approaches consider language and visual cues Pejsa et al. (2014). Human-LLM grounding is crucial for tasks including goal-coordination Bara et al. (2021); Mohanty et al. (2023); Fried et al. (2022); Li and Boyer (2015), planning Chu-Carroll and Carberry (1998); Lochbaum (1998), games Madureira and Schlangen (2023b); Shaikh et al. (2023), data retrieval Lu et al. (2023), improvisation Cho and May (2020), and design Vaithilingam et al. (2024). AI-initiated grounding improves conversation quality Zhou et al. (2022), enables human-AI collaboration Lin et al. (2023), and encourages humans to reflect on LLM outputs Park and Kulkarni (2023). Our work extends this by introducing methods and a benchmark for studying grounding in real-world human-LLM dialogue.
@@
Research has explored the use of machine-learned models to predict and mitigate grounding failures in spoken dialogue systems. By forecasting potential failures, models can guide proactive interventions. For example, in a call-routing system, predictions of downstream grounding issues were used to trigger early transfers to human operators, reducing user frustration and disengagement, like pressing keys to access human assistants or abandoning calls Horvitz and Paek (2007).
@@
| Attempting to advance grounding with the next relevant turn. | ||
| Note: all other successful acts are a subset of Next Turn. | ||
@@ 3 Human-LLM Grounding Acts
To measure grounding between people and LLMs, we curate a set of dialogue acts that serve as proxies for grounding. We outline our selected acts and discuss prior work motivating each act. Our typology builds on prior work in conversational grounding: Clark and Schaefer (1989) and Traum and Hinkelman (1992) outline a hierarchy of actions, including discourse acts, that are used by humans to ground with one another. Recent work has revisited conversational grounding in the context of LLMs Shaikh et al. (2024); Schneider et al. (2024), focusing on a subset of acts generated mainly by people (e.g, following-up, acknowledging understanding, and clarifying).
In contrast, we consider acts generated by both LLMs and people. Our selected acts serve as signals for effective grounding; we segment acts across communicated grounding outcomes. We focus on the construction of common ground, a potential grounding failure, or . Using our typology, we can measure grounding outcomes with observable dialogue acts during human-LLM interaction (illustrative examples in Table 1).
@@
Addressing acts are made in response to detection of inadequate grounding. They explicitly signal a potential misunderstanding. Here, participants engage with a focus on addressing the failure.
Reformulations occur when a participant repeats or restates their query in other words because of a failure to ground. An utterance is a reformulation if the succeeding utterance from the same participant is semantically equivalent to the original. Reformulations are prevalent in search engine and information retrieval domains Lau and Horvitz (1999).
Repairs also signal a grounding failure. Unlike reformulations, the listener directly corrects a misunderstanding from another speaker (e.g. I meant do it in JavaScript, not Python.) Schegloff (1992); Schegloff et al. (1977)
Restarts occur when users reset a conversation to improve understanding and achieve a successful dialogue. They often follow significant misunderstandings, whether in the initial response or across multiple exchanges. Users may restart due to LLMs misinterpreting intent, ambiguity, sensitivity, or irrelevant context Shi et al. (2023), akin to search query retries after irrelevant results Lau and Horvitz (1999). Research on restarts includes user decisions to suspend a dialogue and seek alternate solutions, such as transferring from AI-based systems to human operators before frustration escalates Horvitz and Paek (2007). Restarts are classified when an initial instruction is repaired or reformulated within 30 minutes Radlinski and Joachims (2005); Downey et al. (2007).
@@
The status of grounding may also be uncertain between participants. Rather than clearly indicating success or failure, disambiguating acts represent strategies that participants use to—potentially inefficiently—lower the likelihood of potential misunderstandings.
Clarifications occur when a participant seeks to disambiguate an utterance from another participant; or when a participant proactively “clears up” misunderstandings. Clarifications often occur when the task at hand is perceived as ambiguous (e.g. What did you mean by that?). Ginzburg and Cooper (2001); Purver et al. (2003b, a); Healey et al. (2011, 2003); Purver (2004); Stoyanchev et al. (2013); Madureira and Schlangen (2023a); Rahmani et al. (2023).
Overresponses include more than what another participant reasonably asked for. Unlike Next Turn, which provides only expected information, overresponses also anticipate and respond to potential follow-ups, flouting the Gricean maxim of quantity Grice (1975). Overresponses often appear as overly verbose LLM-generated answers—a behavior contemporary reward models are criticized for encouraging by favoring longer responses Singhal et al. (2023).
@@
Advancing acts signal that a participant understands utterances from another participant.
Next Turns refer to the next conversational move made by a listener that is expected, given the prior turn(s) in a dialogue. Examples of relevant next turns include directly answering a question, expressing an opinion (agreeing or disagreeing), or apologizing. If no misunderstanding has occurred, a listener has moved on to the next relevant turn by default. Levinson et al. (1983); Sacks et al. (1978); Schiffrin (1987). Note that advancing grounding will initiate the Next Relevant Turn. We focus on two: follow-ups & acknowledgments.
Follow-ups elaborate on a prior utterance in an interaction. Unlike clarifications—which disambiguate or clarify—follow-ups seek additional information. Because follow-ups build on a prior utterance, they implicitly signal understanding of past utterances. Davis (1982); Graesser et al. (1995); Traum and Hinkelman (1992); Bunt et al. (2017).
Acknowledgments explicitly signal understanding. These requests manifest either through explicit dialogue (e.g. I see; I understand; O.K.) or by repeating portions of another participant’s utterance (e.g. I can help you [write a story].) Unlike reformulation, where a listener repeats to address failure, acknowledgment occurs when a speaker repeats to demonstrate understanding. Schegloff (1982); Sacks et al. (1978); Schiffrin (1987); Clark and Schaefer (1989); Cho and May (2020)
@@ 4 Data
To analyze collaborative grounding with LLM-based assistants, we draw from three English-language datasets consisting of dialogues between a human and an assistant. WildChat is a real-world human-AI interaction dataset with interaction between people and several OpenAI models Zhao et al. (2024). User data was collected with consent, in exchange for free access to the models. We use the non-toxic version of WildChat, filter conversations to be in English, and sample one conversation from each user. Bing Chat, similar to WildChat, was collected from a large chat-based service used by millions of users, powered by OpenAI LLMs. Finally, MultiWOZ is a crowdsourced dataset of dialogue-based interaction with an assistant Budzianowski et al. (2018); Ramadan et al. (2018); Eric et al. (2019); Zang et al. (2020). In contrast to WildChat and Bing Chat, MultiWOZ contains human-human dialogues, with one human playing a “wizard-of-oz” role as the assistant. We use MultiWOZ 2.2 Zang et al. (2020), examining collaborative grounding acts on the validation and test splits for a subset of the tasks. While we observe similarities in terms of the tasks posed by humans across the three datasets, differences do exist, which makes direct comparisons difficult. In Appendix A.1, we outline the number of dialogues and messages in each dataset.
@@ 5.1 Classifying Human-LLM Grounding Acts
Method.
We employ GPT-4o-mini to annotate grounding acts across a subset of our datasets. On subsets of our data, we observed nearly identical results using GPT-4o compared to GPT-4o-mini. Thus, we employ GPT-4o-mini for its affordability to allow for efficiency and reproducibility. Following Shaikh et al. (2024), we first encode our typology in a prompt (Appendix E.1) and label each turn in the conversations. To validate the accuracy of the approach, three authors annotated 10 dialogues (total of 108 messages) from each dataset. We found that a great deal of the early disagreement among annotators was in assessing clarification vs. follow-up questions. Disgreements were resolved through a round of discussion and reannotation, reaching an average Cohen Kappa of 0.71 across the datasets. For the final dataset, ties were broken through discussion, converging on a final selection of the majority label. Using this as a withheld test set, we find that Macro F-1 scores are reasonably high across datasets (0.75). A full table of results across labels is also in Appendix E.1.
Results.
@@
Before a human ends up addressing grounding, an ideal LLM assistant would have proactively engaged in clarification. However, we find that LLMs clarify at significantly lower rates (, t-test) compared to humans repairing. In fact, the opposite occurs: people clarify LLM outputs ( ) 3 times as much as LLMs clarify user instructions ( ; Fig. 3). In contrast, humans and wizard-of-oz-ed assistants clarify at similar rates (MultiWOZ; human user versus human assistant).
@@
Early grounding patterns also have rippling effects: we find evidence of compounding advancing and addressing patterns as a conversation progresses (Fig. 5). Let denote the message at turn , which can be an advancing grounding act or addressing where and denote the sets of advancing and addressing grounding acts. The likelihood of an utterance representing an addressing act appearing in the first turn on WildChat, and on Bing Chat. We observe a compounding of addressing grounding in human-assistant conversations. If we assume that the turn before was addressing, we find that the likelihood of another failed interaction triples: for WildChat and for Bing Chat. A similar effect appears for advancing acts: early signs of effective conversational grounding similarly snowballs, with on WildChat and on Bing Chat.
5.2 Forecasting Users’ Grounding Acts
In light of the compounding effect of grounding failures (Fig. 5), we pursue the possibility of lowering the probability of grounding failures byintervening before a failure occurs. So far, we have introduced a GPT-based annotator (from §5.1) that annotates conversations with grounding acts in a post-hoc fashion. However, our prompted annotator only identifies opportunities for grounding after they happen. Identifying or curating tasks where grounding failures emerge requires predicting the likelihood of future grounding patterns. To this end, we train a forecaster that predicts grounding acts in the next turn. We focus on WildChat as it is the only publicly availible human-LLM interaction dataset in our analysis.
| Category | Prompt from Rifts Benchmark |